OpenStudy (anonymous):

Instructions:Select the correct answer from each drop-down menu. Find the sum of 2b/(b+a) and 2a\(b^2-a^2 .

OpenStudy (anonymous):

help

OpenStudy (anonymous):

\[\frac{2b}{b+a}+\frac{2a}{b^2-a^2}\]?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

lcd is \(b^2-a^2\)

OpenStudy (anonymous):

since \(b^2-a^2=(b+a)(b-a)\)

OpenStudy (anonymous):

@IrishBoy123